OpenStudy (anonymous):

*ball 2 with with a mass of 30 kg is moving to the left at 10 m/s with that velocity should ball 1 with a mass of 10 kg move to the right and collide with 2 so that ball 1 rebound with a velocity of 30 m/s and ball 2 with velocity of 10 m/s after the collision assume the collision to be perfectly elastic. "momentum" SHOW WORK PLZ

OpenStudy (aravindg):

what we need to find here?

OpenStudy (anonymous):

velocity of ball 1

OpenStudy (aravindg):

do u knw formula for finding this ?

OpenStudy (anonymous):

yes but I plug it in wrong it confuses me a lot.

OpenStudy (amistre64):

you have before and after speeds given, and the masses dont change; so the mvs before have to equal the mvs after right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

m1v1 + m2v2 = m1V1 + m2V2 m1 = 10 kg v1 = V1 = -30 m/s m2 = 30 kg v2 = -10 m/s V2 = 10 m/s

OpenStudy (amistre64):

10v + 30(-10) = 10(-30) + 30(10) 10v -300 = -300 + 300 10v = 300 ifi havent misread it right?

OpenStudy (anonymous):

300/10 = 30 so the velocity is 10 thankyou very much! :D

OpenStudy (amistre64):

hope it worked :)